# scCDCG: Efficient Deep Structural Clustering for single-cell RNA-seq via Deep Cut-informed Graph Embedding

### v1

@@ Abstract
Keywords:
Bioinformatics scRNA-seq data Clustering Deep cut-informed graph embedding Self-supervised learning.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) data analysis is a groundbreaking advancement in the field of bioinformatics that enables researchers to explore the complex structure and heterogeneity within cell populations and reveal different cell types, states, or subpopulations [18]. Clustering is an important tool for analyzing scRNA-seq data, which separates the cells into different groups with similar cells in the same group, thus can reveal the complex characteristics of different cell populations [12]. Clustering of scRNA-seq data has proven to be particularly valuable in elucidating cell development, understanding disease mechanisms, and identifying novel cell types and subtypes [8].
@@
To bridge this gap, a series of methods that create cell-cell graphs based on relationships between cells have emerged to incorporate intercellular structural information [7]. Moreover, in order to introduce high-order structural information, the latest scRNA-seq data clustering methods have started to use Graph Neural Networks (GNNs). Specifically, scGNN [27] utilizes GNNs and three multi-modal autoencoders to formulate and aggregate cell-cell relationships; scDSC [6] combines a ZINB-based autoencoder and GNNs in a mutually supervised manner, enhancing the clustering performance. Despite considering intercellular structure information in scRNA-seq data, the GNN-based clustering still has some shortcomings: (i) As single-cell transcriptome sequencing technology continues to advance, the scale of scRNA-seq data is getting increasingly larger. However, GNNs have high complexity and poor efficiency, i.e., longer training/inference time, especially when facing large-scale datasets [29]. (ii) GNNs tend to generate similar representations for all nodes [17, 32, 28], which will even lead to the over-smoothing issue [4] (i.e., indistinguishable representations of nodes in different classes) when GNNs become deeper (i.e., more layers). This problem will become more severe when GNNs face the scRNA-seq data, because scRNA-seq data’s high-dimension and high-sparsity will result in the similarity between all cell being close to 0, thus all cells are more poorly distinguishable. Therefore, GNN-based clustering methods face challenges in handling the inefficiency due to scRNA-seq data’s intrinsic characteristics.
@@
Thus, it is crucial to develop a scRNA-seq data clustering method that simultaneously utilizes intercellular high-order structural information, handles high-dimensional and high-sparse scRNA-seq datasets, and demonstrates high efficiency. In order to achieve this goal, we introduce scCDCG (single-cell RNA-seq Clustering via Deep Cut-informed Graph). The scCDCG consists of three main modules, specifically: (i) scCDCG incorporates complex intercellular high-order structural information through a novel and simple method of graph embedding module based on deep cut-informed techniques [19], while avoiding the efficiency limitations and over-smoothing issue of GNN-based methods. It is well-suited for high-dimensional and high-sparse scRNA-seq data, demonstrating enhanced efficiency. (ii) scCDCG utilizes a new self-supervised learning module based on optimal transport to guide the model learning process [15], which is more adaptive to high-dimensional and high-sparse scRNA-seq data and enhances clustering accuracy and robustness. (iii) scCDCG employs an autoencoder-based feature learning module for dimension reduction and feature extraction, effectively reducing the model complexity. To fully verify the capacity and effectiveness of scCDCG, we conduct extensive experiments with 7 established models on 6 datasets, the results demonstrate the superior clustering performance and efficiency of scCDCG. In addition, we conduct sufficient ablation and analysis experiments to validate the effectiveness of each module of scCDCG.
@@ 3.2 The framework of scCDCG
$\arg\min_{\bm{H}}\text{Tr}(\bm{H}^{T}L_{\bm{C}}\bm{H})$ subject to $\text{subject to }\bm{H}^{T}\bm{H}=\bm{I},$ (6)
$\arg\min_{\bm{H}}\text{Tr}(\bm{H}^{T}L_{\bm{S}}\bm{H})$ subject to $\text{subject to }\bm{H}^{T}\bm{H}=\bm{I},$
@@
After obtaining the clustering result distribution $Q$, we aim to optimize the data representation by learning the high-confidence assignments. This approach aims to minimize the difference between the data representations and the cluster centers, thereby improving the cohesion of the clusters. For instance, SDCN [2] calculated a target distribution $P=\left[p_{ij}\right]$:
@@
Here, we consider the target distribution $P$ as the transport plan matrix within the optimal transport theory and $-\text{log}Q$ as the associated cost matrix. The constraint $P^{T}\bm{1}_{N}=N\bm{\pi}$ is imposed, with $\pi$ representing the proportion of points in each cluster, which can be estimated by the intermediate clustering result. This step successfully adds the constraint that the result cluster distribution must be consistent with the mixing proportions. Given the computational burden of direct optimization, we leverage the Sinkhorn distance [20] for rapid optimization through an entropic constraint. The optimization problem with an entropy constraint Lagrange multiplier is as follows:
$\min_{P}\quad$ $-P*(\text{log}Q)-\frac{1}{\lambda}\text{H}(P)$ (12)
@@
$\vspace{-2mm}\bm{T}^{*}=\text{diag}(\bm{u}^{(t)})Q^{\lambda}\text{diag}(\bm{v}^{(t)}),$ (13)
@@ 4.3 Ablation Study
Table 3: Ablation study of orthogonality regularization and optimal transport. Dataset Metric scCDCG -w/o Ort scCDCG -w/o OT scCDCG CITE-CMBC ACC 56.82 69.45 71.45 NMI 57.99 71.83 79.34 ARI 45.47 81.26 81.26 Worm Neuron cells ACC 63.16 70.18 78.73 NMI 53.82 65.35 72.11 ARI 40.43 50.88 59.78 Mouse Bladder cells ACC 62.97 70.89 75.61 NMI 69.05 65.21 75.91 ARI 54.55 58.45 64.55 Human Pan -creas cells 2 ACC 58.77 70.03 84.66 NMI 68.14 70.87 83.44 ARI 55.31 62.41 85.30

### v2

@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) data analysis is a groundbreaking advancement in the field of bioinformatics that enables researchers to explore the complex structure and heterogeneity within cell populations and reveal different cell types, states, or subpopulations [21, 31, 36]. Clustering is an important tool for analyzing scRNA-seq data, which separates the cells into different groups with similar cells in the same group, thus can reveal the complex characteristics of different cell populations [11]. Clustering of scRNA-seq data has proven to be particularly valuable in elucidating cell development, understanding disease mechanisms, and identifying novel cell types and subtypes [35].
@@
To bridge this gap, a series of methods that create cell-cell graphs based on relationships between cells have emerged to incorporate intercellular structural information [7]. Moreover, in order to introduce high-order structural information, the latest scRNA-seq data clustering methods have started to use Graph Neural Networks (GNNs). Specifically, scGNN [30] utilizes GNNs and three multi-modal autoencoders to formulate and aggregate cell-cell relationships; scDSC [6] combines a ZINB-based autoencoder and GNNs in a mutually supervised manner, enhancing the clustering performance. Despite considering intercellular structure information in scRNA-seq data, the GNN-based clustering still has some shortcomings: (i) As single-cell transcriptome sequencing technology continues to advance, the scale of scRNA-seq data is getting increasingly larger. However, GNNs have high complexity and poor efficiency, i.e., longer training/inference time, especially when facing large-scale datasets [33]. (ii) GNNs tend to generate similar representations for all nodes [19, 38, 32, 18], which will even lead to the over-smoothing issue [4] (i.e., indistinguishable representations of nodes in different classes) when GNNs become deeper (i.e., more layers). This problem will become more severe when GNNs face the scRNA-seq data, because scRNA-seq data’s high-dimension and high-sparsity will result in the similarity between all cell being close to 0, thus all cells are more poorly distinguishable. Therefore, GNN-based clustering methods face challenges in handling the inefficiency due to scRNA-seq data’s intrinsic characteristics.
@@
Thus, it is crucial to develop a scRNA-seq data clustering method that simultaneously utilizes intercellular high-order structural information, handles high-dimensional and high-sparse scRNA-seq datasets, and demonstrates high efficiency. In order to achieve this goal, we introduce scCDCG (single-cell RNA-seq Clustering via Deep Cut-informed Graph). The scCDCG consists of three main modules, specifically: (i) scCDCG incorporates complex intercellular high-order structural information through a novel and simple method of graph embedding module based on deep cut-informed techniques [22, 20], while avoiding the efficiency limitations and over-smoothing issue of GNN-based methods [16]. It is well-suited for high-dimensional and high-sparse scRNA-seq data, demonstrating enhanced efficiency. (ii) scCDCG utilizes a new self-supervised learning module based on optimal transport to guide the model learning process [14], which is more adaptive to high-dimensional and high-sparse scRNA-seq data and enhances clustering accuracy and robustness. (iii) scCDCG employs an autoencoder-based feature learning module for dimension reduction and feature extraction, effectively reducing the model complexity. To fully verify the capacity and effectiveness of scCDCG, we conduct extensive experiments with 7 established models on 6 datasets, the results demonstrate the superior clustering performance and efficiency of scCDCG. In addition, we conduct sufficient ablation and analysis experiments to validate the effectiveness of each module of scCDCG.
@@ 3.2 The framework of scCDCG
$\arg\min_{\bm{H}}\text{Tr}(\bm{H}^{T}L_{\bm{C}}\bm{H})$ $\text{subject to }\bm{H}^{T}\bm{H}=\bm{I},$ (6)
$\arg\min_{\bm{H}}\text{Tr}(\bm{H}^{T}L_{\bm{S}}\bm{H})$ $\text{subject to }\bm{H}^{T}\bm{H}=\bm{I},$
@@
After obtaining the clustering result distribution $Q$, we aim to optimize the data representation by learning the high-confidence assignments [17]. This approach aims to minimize the difference between the data representations and the cluster centers, thereby improving the cohesion of the clusters. For instance, SDCN [2] calculated a target distribution $P=\left[p_{ij}\right]$:
@@
Here, we consider the target distribution $P$ as the transport plan matrix within the optimal transport theory and $\text{log}Q$ as the associated cost matrix. The constraint $P^{T}\bm{1}_{N}=N\bm{\pi}$ is imposed, with $\pi$ representing the proportion of points in each cluster, which can be estimated by the intermediate clustering result. This step successfully adds the constraint that the result cluster distribution must be consistent with the mixing proportions. Given the computational burden of direct optimization, we leverage the Sinkhorn distance [23] for rapid optimization through an entropic constraint. The optimization problem with an entropy constraint Lagrange multiplier is as follows:
$\min_{P}$ $-P*(\text{log}Q)-\frac{1}{\lambda}\text{H}(P)$ (12)
@@
$\vskip-5.69054pt\hat{P}^{(t)}=\operatorname{diag}\left(\bm{u}^{(t)}\right)Q^{\lambda}\operatorname{diag}\left(\bm{v}^{(t)}\right)$ (13)
@@ 4.3 Ablation Study
Table 3: Ablation study of orthogonality regularization and optimal transport.
Dataset Metric scCDCG -w/o Ort scCDCG -w/o OT scCDCG
CITE-CMBC ACC 56.82 69.45 71.45
NMI 57.99 71.83 79.34
ARI 45.47 81.26 81.26
Worm Neuron cells ACC 63.16 70.18 78.73
NMI 53.82 65.35 72.11
ARI 40.43 50.88 59.78
Mouse Bladder cells ACC 62.97 70.89 75.61
NMI 69.05 65.21 75.91
ARI 54.55 58.45 64.55
Human Pan -creas cells 2 ACC 58.77 70.03 84.66
NMI 68.14 70.87 83.44
ARI 55.31 62.41 85.30